Rogen Corporation manufactures a single product. The standard cost per unit of product is shown below.

Direct materials—1 pound plastic at $6.00 per pound $ 6.00
Direct labor—2.0 hours at $11.85 per hour 23.70
Variable manufacturing overhead 12.00
Fixed manufacturing overhead 12.00
Total standard cost per unit $53.70


The predetermined manufacturing overhead rate is $12.00 per direct labor hour ($24.00 ÷ 2.0). It was computed from a master manufacturing overhead budget based on normal production of 11,800 direct labor hours (5,900 units) for the month. The master budget showed total variable costs of $70,800 ($6.00 per hour) and total fixed overhead costs of $70,800 ($6.00 per hour). Actual costs for October in producing 3,700 units were as follows.

Direct materials (3,890 pounds) $ 23,729
Direct labor (7,300 hours) 87,965
Variable overhead 63,934
Fixed overhead 26,966
    Total manufacturing costs $202,594


The purchasing department buys the quantities of raw materials that are expected to be used in production each month. Raw materials inventories, therefore, can be ignored.

(a)

Compute all of the materials and labor variances.

Total materials variance $

UnfavorableFavorableNeither favorable nor unfavorable

Materials price variance $

Neither favorable nor unfavorableFavorableUnfavorable

Materials quantity variance $

UnfavorableFavorableNeither favorable nor unfavorable

Total labor variance $

Neither favorable nor unfavorableFavorableUnfavorable

Labor price variance $

FavorableNeither favorable nor unfavorableUnfavorable

Labor quantity variance $

FavorableUnfavorableNeither favorable nor unfavorable


(b)

Compute the total overhead variance.

Total overhead variance $

UnfavorableFavorableNeither favorable nor unfavorable
Answer #1
Total Material Variance = Standard cost - Actual Cost
(6X3700)-23729=$1529 UF
calculation of direct material price variance:
= (Standard price per unit of material - Actual price per unit of material) × Actual quantity
= ($6/Pound - $6.10/Pound ) × 3890 Pound = $389 U F
Calculation of direct material quantity variance
=(standard quantity of material required for actual production - actual quantity used) × Standard price per unit
((3700-3890 Pound )X $6 = $1140 UF
* Actual Price/ Pound= 23729/3890=$6.10
Total Labour Variance= Stdandard Cost- Actual Cost
(23.70X 3700)-87965=$275 UF
Calculation of direct labor rate variance
= (Standard direct labor rate per hour - actual direct labor rate per hour) × Actual hours used
= ($11.85/hour - $12.05/hour) × 7300 Hours= $1460 UF
* Actual Hourly Rate:=$87965/7300=$12.05/Hour
Calculation of direct labor efficiency variance:
= (standard hours required for actual production - actual hours used) × standard overhead recovery rate
= (2 Hour × 3700 Unit - 7300) × $11.85 = $1185 F
Total Overhead Variance= Applied Overhead- Actual Overhead
($12X (2 HourX 3700 Unit)- (63934+26966)
(88800-90900)= $2100 UF